# Fixed-Time Input-to-State Stability for Singularly Perturbed Systems via Composite Lyapunov Functions

### v1

@@ 3.1 Main Results
where $V,W,a_{i},b_{i},k_{i},\kappa_{i}$ come from Assumptions 2-3. With these terms at hand, we can now state the first main result of the paper.
@@
Consider system (7), and suppose Assumptions 1-3 hold. Furthermore, suppose there exists $\nu_{1},\nu_{2},\omega_{1},\omega_{2}\in\mathbb{R}$ and $\rho_{1},\rho_{2}\in\mathcal{K}_{\infty}$ such that the terms in (9) satisfy
$I_{1}$ $\leq\nu_{1}\tilde{V}^{2}+\omega_{1}\tilde{W}^{2}+\rho_{1}(|u|)$ (11a)
$I_{2}$ $\leq\nu_{2}\tilde{V}^{2}+\omega_{2}\tilde{W}^{2}+\rho_{2}(|u|)$ (11b)
@@
Then there exists $\varepsilon^{*}$ such that (7) is FxT ISS for $\varepsilon\in(0,\varepsilon^{*})$.
@@
Consider the following composite Lyapunov function candidate
$\Psi_{\zeta}(x,y)=\zeta V(x)+(1-\zeta)W(x,y),\quad\zeta\in(0,1)$ (12)
where $V$ and $W$ are from Assumptions 2 and 3 respectively. By denoting $m_{\zeta}=\min(\zeta,1-\zeta)$, $\varrho_{1}=\min(\alpha_{1},\beta_{1})$ and $\varrho_{2}=\max(\alpha_{2},\beta_{2})$, it can be verified that for each fixed $\zeta\in(0,1)$, $\Psi_{\zeta}$ satisfies the following:
$m_{\zeta}\varrho_{1}\left(\frac{1}{2}\left\lvert\begin{bmatrix}x\\ y\end{bmatrix}\right\rvert\right)\leq\Psi_{\zeta}(x,y)\leq 2(1-m_{\zeta})\varrho_{2}\left(\left\lvert\begin{bmatrix}x\\ y\end{bmatrix}\right\rvert\right).$
Now we can compute the Lie derivative of (12) along the trajectories of (7):
$\dot{\Psi_{\zeta}}$ $=\zeta\left(\frac{\partial V}{\partial x}f(x,h(x),u)+I_{1}\right)$
$~{}~{}~{}+(1-\zeta)\left(\frac{1}{\varepsilon}\frac{\partial W}{\partial y}g(x,y+h(x),u)+I_{2}\right).$
We can use Assumptions 2 and 3 to obtain:
$\dot{\Psi_{\zeta}}$ $\leq-\zeta\underline{k}(V^{a_{1}}+V^{a_{2}})+\zeta\rho_{R}(|u|)+\zeta I_{1}$
$~{}~{}~{}-\frac{1-\zeta}{\varepsilon}\underline{\kappa}(W^{b_{1}}+W^{b_{2}})+\frac{1-\zeta}{\varepsilon}\rho_{B}(|u|)+(1-\zeta)I_{2}$
$\leq-\dfrac{\zeta}{2}\underline{k}\tilde{V}^{2}-\dfrac{1-\zeta}{2\varepsilon}\underline{\kappa}\tilde{W}^{2}+\zeta I_{1}+(1-\zeta)I_{2}$
$~{}~{}~{}+\zeta\rho_{R}(|u|)+\frac{1-\zeta}{\varepsilon}\rho_{B}(|u|).$
With conditions (11a) and (11b), we have:
$\dot{\Psi_{\zeta}}$ $\leq-\nu(\zeta)\tilde{V}^{2}-\omega_{\varepsilon}(\zeta)\tilde{W}^{2}+\rho_{\varepsilon,\zeta}(|u|),$
@@
$\rho_{\varepsilon,\zeta}(|u|)$ $=\zeta\left(\rho_{R}(|u|)+\rho_{1}(|u|)\right)$
$~{}~{}~{}+(1-\zeta)\left(\frac{\rho_{B}(|u|)}{\varepsilon}+\rho_{2}(|u|)\right)$
and we clearly have $\rho_{\varepsilon,\zeta}\in\mathcal{K}_{\infty}$ for $\zeta\in(0,1)$ and $\varepsilon>0$. By (11c) we can find $\zeta^{*}\in(0,1)$ such that $\nu(\zeta^{*})>0$, and then we can find $\varepsilon^{*}>0$ such that $\omega_{\varepsilon}(\zeta^{*})>0$ whenever $\varepsilon\in(0,\varepsilon^{*})$. For these $\varepsilon\in(0,\varepsilon^{*})$ we can let $\lambda_{\varepsilon}=\min\left(\nu(\zeta^{*}),\omega_{\varepsilon}(\zeta^{*})\right)>0$ and obtain:
$\dot{\Psi}_{\zeta^{*}}$ $\leq-\lambda_{\varepsilon}\left(\tilde{V}^{2}+\tilde{W}^{2}\right)+\rho_{\varepsilon,\zeta}(|u|)$
$\leq-\lambda_{\varepsilon}\left(V^{a_{1}}+V^{a_{2}}+W^{b_{1}}+W^{b_{2}}\right)+\rho_{\varepsilon,\zeta}(|u|)$
$=-\dfrac{\lambda_{\varepsilon}}{2}\bigl{(}(V^{a_{1}}+V^{a_{2}})+(V^{a_{1}}+V^{a_{2}})+(W^{b_{1}}+W^{b_{2}})$
$~{}~{}~{}+(W^{b_{1}}+W^{b_{2}})\bigl{)}+\rho_{\varepsilon,\zeta}(|u|).$
We can pick $\gamma_{1}\in\left[\max\left(a_{1},b_{1}\right),1\right)$ and $\gamma_{2}\in\left(1,\min\left(a_{2},b_{2}\right)\right]$ to obtain
$\dot{\Psi}_{\zeta^{*}}$ $\leq-\dfrac{\lambda_{\varepsilon}}{2}\left(V^{\gamma_{1}}+V^{\gamma_{2}}+W^{\gamma_{1}}+W^{\gamma_{2}}\right)+\rho_{\varepsilon,\zeta}(|u|)$
$\leq-\dfrac{\lambda_{\varepsilon}}{2}\left((V+W)^{\gamma_{1}}+2^{1-\gamma_{2}}(V+W)^{\gamma_{2}}\right)+\rho_{\varepsilon,\zeta}(|u|)$
$\leq-\dfrac{\lambda_{\varepsilon}}{2}\left(\Psi_{\zeta^{*}}^{\gamma_{1}}+2^{1-\gamma_{2}}\Psi_{\zeta^{*}}^{\gamma_{2}}\right)+\rho_{\varepsilon,\zeta}(|u|).$
@@
The case where the right hand sides of (11a) and (11b) also contain a $c_{i}\tilde{V}\tilde{W}$ term can be reduced to (11a) and (11b) by applying Lemma 1. $\square$
We have shown that under suitable assumptions, the inequalities in (11) imply that system (7) is FxT ISS provided there is a sufficiently large time scale separation. However, the question of whether (1) is FxT ISS may also be of interest. In other words: we ask if the transformation $y=z-h(x)$ preserves the FxT ISS property. Fortunately, as long as the quasi steady state map satisfies a mild boundedness assumption, (7) being FxT ISS implies (1) is FxT ISS. This is further detailed in the following result.
@@
Suppose (7) is FxT ISS and $|h(x)|\leq\tilde{\alpha}(|x|)$ for some $\tilde{\alpha}\in\mathcal{K}$; then (1) is also fixed time ISS. Furthermore, if $\Psi(x,y)$ is a FxT ISS Lyapunov function for (7), then $\tilde{\Psi}(x,z):=\Psi(x,z-h(x))$ is a FxT ISS Lyapunov function for (1).
Proof 3.3.
Since (7) is FxT ISS, there exists $\beta\in\mathcal{GKL}$ and $\vartheta\in\mathcal{K}$ such that
$\left\lvert\begin{bmatrix}x(t)\\ z(t)-h(x(t))\end{bmatrix}\right\rvert\leq\beta\left(\left\lvert\begin{bmatrix}x(0)\\ y(0)\end{bmatrix}\right\rvert,t\right)+\vartheta(|u|_{\infty}),$ (14)
where $\beta(r,t)=0$ for each $t>{T}_{\beta}(r)$, and ${T}_{\beta}$ is a continuous function that satisfies $\sup_{r\geq 0}{T}_{\beta}(r)<\tilde{T}<\infty$ and $T_{\beta}(0)=0$. Without loss of generality, we can assume $\beta(\cdot,t)$ is non-decreasing for each $t\geq 0$. Note that (14) implies
$|x(t)|$ $\leq\beta\left(\left\lvert\begin{bmatrix}x(0)\\ y(0)\end{bmatrix}\right\rvert,t\right)+\vartheta(|u|_{\infty})$ (15a)
$|z(t)|$ $\leq\beta\left(\left\lvert\begin{bmatrix}x(0)\\ y(0)\end{bmatrix}\right\rvert,t\right)+\vartheta(|u|_{\infty})+|h(x(t))|.$ (15b)
@@
$\beta\left(\left\lvert\begin{bmatrix}x(0)\\ y(0)\end{bmatrix}\right\rvert,t\right)$ $=\beta\left(\left\lvert\begin{bmatrix}x(0)\\ z(0)-h(x(0))\end{bmatrix}\right\rvert,t\right)$
$\leq\beta\left(2\left\lvert\begin{bmatrix}x(0)\\ z(0)\end{bmatrix}\right\rvert,t\right)+\beta\left(2\tilde{\alpha}(|x(0)|),t\right)$
$\leq\overline{\beta}\left(\left\lvert\begin{bmatrix}x(0)\\ z(0)\end{bmatrix}\right\rvert,t\right),$
where $\overline{\beta}(r,t):=\beta(2r,t)+\beta(2\tilde{\alpha}(r),t)\in\mathcal{GKL}$. We also have
$|h(x(t))|$ $\leq\tilde{\alpha}(|x(t)|)$
$\leq\tilde{\alpha}\left(\overline{\beta}\left(\left\lvert\begin{bmatrix}x(0)\\ z(0)\end{bmatrix}\right\rvert,t\right)+{\vartheta}(|u|_{\infty})\right)$
$\leq\tilde{\beta}\left(\left\lvert\begin{bmatrix}x(0)\\ z(0)\end{bmatrix}\right\rvert,t\right)+\tilde{\vartheta}(|u_{\infty}|),$ (16)
where $\tilde{\beta}(r,t)=\tilde{\alpha}\left(2\overline{\beta}\left(r,t\right)\right)\in\mathcal{GKL}$, and $\tilde{\vartheta}(\cdot)=\tilde{\alpha}(2{\vartheta}(\cdot))\in\mathcal{K}$. Combining (15) and (16) yields first result.
Now suppose $\Psi(x,y)$ is a FxT ISS Lyapunov function for (7). By the hypothesis we already know $\tilde{\Psi}(x,z)$ will satisfy (5) along system (1), so we only need to check (4). Indeed, since $\Psi$ satisfies (4) we have $\tilde{\Psi}(x,z)\leq\alpha_{2}^{*}\left(\left\lvert[x,z]^{\top}\right\rvert\right)$, where $\alpha_{2}^{*}(r):=\alpha_{2}(2r)+\alpha_{2}(2\tilde{\alpha}(r))\in\mathcal{K}_{\infty}$. Now we pick $\alpha_{1}^{*}\in\mathcal{K}_{\infty}$ satisfying
$\alpha_{1}^{*}(2r)+\alpha_{1}^{*}(2\tilde{\alpha}(r))\leq\alpha_{1}(r),\quad r\geq 0$
To see why such $\alpha_{1}^{*}$ exists, we first want $\alpha_{1}^{*}(2r)\leq\frac{1}{2}\alpha_{1}(r)$, which is accomplished when $\alpha_{1}^{*}(r)\leq\frac{1}{2}\alpha_{1}(\frac{1}{2}r)\in\mathcal{K}_{\infty}$. We pick $\overline{\alpha}\in\mathcal{K}_{\infty}$ satisfying $\overline{\alpha}(r)\geq\tilde{\alpha}_{1}(r)$ for all $r\geq 0$. We want $\alpha_{1}^{*}(2\tilde{\alpha}_{1}(r))\leq\frac{1}{2}\alpha_{1}(r)$, which is satisfied whenever $(\alpha_{1}^{*})^{-1}(\frac{1}{2}\alpha_{1}(r))\geq 2\overline{\alpha}(r)\Leftrightarrow(\alpha_{1}^{*})^{-1}(r)\geq 2\overline{\alpha}(\alpha_{1}^{-1}(2r))\Leftrightarrow\alpha_{1}^{*}(r)\leq\alpha^{-1}(r)$, where $\alpha=2\overline{\alpha}\circ\alpha_{1}^{-1}\circ 2\text{Id}\in\mathcal{K}_{\infty}$ (here Id denotes the identity function), so we can set $\alpha_{1}^{*}(r)=\min(\frac{1}{2}\alpha_{1}(\frac{1}{2}r),\alpha^{-1}(r))\in\mathcal{K}_{\infty}$. Then, we have $\alpha_{1}^{*}\left(\left\lvert[x,z]^{\top}\right\rvert\right)\leq\alpha_{1}\left(\left\lvert[x,y]^{\top}\right\rvert\right)$, which concludes the proof.
@@ 3.2 A Stylized Example
To illustrate our results, we first consider a simple interconnection between two scalar systems. Consider the plant
@@
where $\lceil\cdot\rfloor^{q}:=|\cdot|^{q}\text{sgn}(\cdot)$, $x,z,u_{1},u_{2}\in\mathbb{R}$, $10<q_{1}\leq r_{1}<$ 1 and $1<r_{2}\leq q_{2}$. It is easy to check that the system has quasi steady state $h(x)=x$ with reduced system:
@@
and boundary layer system:

### v2

@@ 3.2 Analysis
where $V,W,a_{i},b_{i},k_{i},\kappa_{i}$ come from Assumptions 2-3.
With these definitions at hand, we can now state the first main result of the paper.
@@
Consider system (9), and suppose that Assumptions 1-3 hold. Furthermore, suppose there exists $\nu_{1},\nu_{2},\omega_{1},\omega_{2}\in\mathbb{R}$ and $\rho_{1},\rho_{2}\in\mathcal{K}_{\infty}$ such that the interconnection terms in (13) satisfy
$I_{1}{(x,y,u)}$ $\leq\nu_{1}\tilde{V}^{2}{(x)}+\omega_{1}\tilde{W}^{2}{(x,y)}+\rho_{1}(|u|)$ (15a)
$I_{2}{(x,y,u)}$ $\leq\nu_{2}\tilde{V}^{2}{(x)}+\omega_{2}\tilde{W}^{2}{(x,y)}+\rho_{2}(|u|)$ (15b)
@@
Then, there exists $\varepsilon^{*}>0$ such that, for each $\varepsilon\in(0,\varepsilon^{*})$, the system (9) is FxT ISS. $\square$
@@
Continuing from (12), we can use Assumptions 2 and 3 to obtain:
$\dot{\Psi}_{\zeta}$ $\leq-\zeta\underline{k}(V^{a_{1}}{(x)}+V^{a_{2}}{(x)})+\zeta\rho_{R}(|u|)+\zeta I_{1}+(1-\zeta)I_{2}$
$~~~-\frac{1-\zeta}{\varepsilon}\underline{\kappa}(W^{b_{1}}{(x,y)}+W^{b_{2}}{(x,y)})+\frac{1-\zeta}{\varepsilon}\rho_{B}(|u|)$
$\leq-\dfrac{\zeta}{2}\underline{k}\tilde{V}^{2}{(x)}-\dfrac{1-\zeta}{2\varepsilon}\underline{\kappa}\tilde{W}^{2}{(x,y)}+\zeta I_{1}+(1-\zeta)I_{2}$
$~~~+\zeta\rho_{R}(|u|)+\frac{1-\zeta}{\varepsilon}\rho_{B}(|u|),$
where, for simplicity, we omit the arguments of $I_{1}$ and $I_{2}$. With conditions (15a) and (15b), we have:
$\dot{\Psi}_{\zeta}$ $\leq-\nu(\zeta)\tilde{V}^{2}{(x)}-\omega_{\varepsilon}(\zeta)\tilde{W}^{2}{(x,y)}+\rho_{\varepsilon,\zeta}(|u|),$
@@
$\rho_{\varepsilon,\zeta}(s)$ $=\zeta\left(\rho_{R}(s)+\rho_{1}(s)\right)+(1-\zeta)\left(\frac{\rho_{B}(s)}{\varepsilon}+\rho_{2}(s)\right),$
and we clearly have $\rho_{\varepsilon,\zeta}\in\mathcal{K}_{\infty}$ for $\zeta\in(0,1)$ and $\varepsilon>0$. By (15c) we can find $\zeta^{*}\in(0,1)$ such that $\nu^{*}:=\nu(\zeta^{*})>0$, and then we can find $\varepsilon^{*}>0$ such that $\omega_{\varepsilon}(\zeta^{*})>\nu^{*}$ whenever $\varepsilon\in(0,\varepsilon^{*})$. For these $\varepsilon\in(0,\varepsilon^{*})$ we obtain:
$\dot{\Psi}_{\zeta^{*}}$ $\leq-\nu^{*}\left(\tilde{V}^{2}{(x)}+\tilde{W}^{2}{(x,y)}\right)+\rho_{\varepsilon,\zeta^{*}}(|u|)$
$\leq-\nu^{*}\left(V^{a_{1}}{(x)}+V^{a_{2}}{(x)}+W^{b_{1}}{(x,y)}+W^{b_{2}}{(x,y)}\right)$
$~~~+\rho_{\varepsilon,\zeta^{*}}(|u|)$
$=-\dfrac{\nu^{*}}{2}\bigl{(}(V^{a_{1}}{(x)}+V^{a_{2}}{(x)})+(V^{a_{1}}{(x)}+V^{a_{2}}{(x)})$
$~~~+(W^{b_{1}}{(x,y)}+W^{b_{2}}{(x,y)})+(W^{b_{1}}{(x,y)}+W^{b_{2}}{(x,y)})\bigl{)}$
$~~~+\rho_{\varepsilon,\zeta^{*}}(|u|).$
We pick $\gamma_{1}\in\left[\max\left\{a_{1},b_{1}\right\},1\right)$ and $\gamma_{2}\in\left(1,\min\left\{a_{2},b_{2}\right\}\right]$ to obtain
$\dot{\Psi}_{\zeta^{*}}$ $\leq-\dfrac{\nu^{*}}{2}\left(V^{\gamma_{1}}{(x)}+V^{\gamma_{2}}{(x)}+W^{\gamma_{1}}{(x,y)}+W^{\gamma_{2}}{(x,y)}\right)$
$~~~+\rho_{\varepsilon,\zeta^{*}}(|u|)$
$\leq-\dfrac{\nu^{*}}{2}\left((V{(x)}+W{(x,y)})^{\gamma_{1}}+2^{1-\gamma_{2}}(V{(x)}+W{(x,y)})^{\gamma_{2}}\right)$
$~~~+\rho_{\varepsilon,\zeta^{*}}(|u|)$
$\leq-\dfrac{\nu^{*}}{2}\left(\Psi_{\zeta^{*}}^{\gamma_{1}}+2^{1-\gamma_{2}}\Psi_{\zeta^{*}}^{\gamma_{2}}\right)+\rho_{\varepsilon,\zeta^{*}}(|u|).$
@@
We would like to note that the “or” condition in (15c) is non-exclusive, i.e it is acceptable for both conditions to be satisfied. In this case, any choice of the weight $\zeta\in(0,1)$ will result in a valid composite Lyapunov function candidate.
@@
The functions $\tilde{V}(x)$ and $\tilde{W}(x,y)$ are highly analogous to the functions $\psi_{1}(y)$ and $\psi_{2}(y)$, respectively, from [7, Chapter 11.5]. For the asymptotic stability (resp. FxT ISS) result from [7, Theorem 11.3] (resp. Theorem 1 in this paper), the Lie derivatives of $V$ and $W$ along the reduced and boundary layer dynamics are assumed to be upper bounded by negative multiples of $\psi_{1}^{2}(x)$ and $\psi_{2}^{2}(y)$ (resp. $\tilde{V}^{2}(x)$ and $\tilde{W}^{2}(x,y)$), respectively. Hence, we can observe that the structure of the interconnection conditions in our paper are actually more forgiving than those from [7, Chapter 11.5] in the sense that we allow for extra $\tilde{V}^{2}(x)$ and $\tilde{W}^{2}(x,y)$ terms. This extra degree of freedom is particularly useful since it allows us to apply our results to a variety of interesting systems, such as those presented later in this paper. However, our conditions are also more restrictive in the sense that we now require very specific forms for the expressions $\psi_{1}$ and $\psi_{2}$. But this is expected, since Theorem 1 considers fixed-time ISS, which is a much stronger notion of stability.
We have shown that under suitable assumptions, the inequalities in (15) imply that system (9) is FxT ISS provided there is a sufficiently large time scale separation. The proof of Theorem 1 provides an efficient methodology for computing a somewhat conservative estimate of the required timescale separation, i.e $\varepsilon^{*}$, for FxT ISS. It can also be seen from the proof of Theorem 1 that if the fast dynamics have no input, then the derived gain $\varrho$ in the FxT ISS bound (4) can be made independent of $\varepsilon$. This is further detailed in the following Corollary.
Corollary 2.
Consider (9) and suppose the conditions of Theorem 1 are satisfied, but with $\rho_{B}\equiv 0$ and the vector field $g$ is independent of $u$. Then, there exists $\beta\in\mathcal{GKL}$, $\varrho\in\mathcal{K}$ and $\varepsilon^{*}>0$ such that the following holds
$\left\lvert s(t)\right\rvert\leq\beta\left(\left\lvert s_{0}\right\rvert,t\right)+\varrho(|u|_{\infty}),$ (17)
for all $t\geq 0$, $u\in\mathcal{L}_{\infty}^{p}$, and $\varepsilon\in(0,\varepsilon^{*})$, where $s(t):=[x(t),y(t)]^{\top}$.
Proof 3.3.
We follow the same steps from the proof of Theorem 1, but with $\tilde{\rho}_{\zeta}(s)=\zeta\left(\rho_{R}(s)+\rho_{1}(s)\right)+(1-\zeta)\rho_{2}(s)$ instead of $\rho_{\varepsilon,\zeta}$. This results in an upper bound on $\dot{\Psi}_{\zeta^{*}}$ that holds uniformly for $\varepsilon\in(0,\varepsilon^{*})$.
While the results of Theorem 1 and Corollary 2 hold for system (9), the question of whether (1) is FxT ISS may also be of interest. In other words: we ask if the transformation y=zh(x)y=z-h(x) preserves the FxT ISS property. Fortunately, as long as the quasi steady state map satisfies a mild boundedness assumption, (9) being FxT ISS implies (1) is FxT ISS. This is further detailed in the following result.
@@
Suppose (9) is FxT ISS and $|h(x)|\leq\tilde{\alpha}(|x|)$ for some $\tilde{\alpha}\in\mathcal{K}$, then (1) is also fixed time ISS. $\square$
Proof 3.4.
Since (9) is FxT ISS, there exists $\beta\in\mathcal{GKL}$ and $\varrho\in\mathcal{K}$ such that
$\left\lvert\begin{bmatrix}x(t)\\ z(t)-h(x(t))\end{bmatrix}\right\rvert\leq\beta\left(\left\lvert\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}\right\rvert,t\right)+\varrho(|u|_{\infty}),$ (18)
where $\beta(r,t)=0$ for each $t>T(r)$, and $T$ is a continuous function that satisfies $\sup_{r\geq 0}T(r)<\infty$ and $T(0)=0$. Without loss of generality, we can assume $\beta(\cdot,t)$ is non-decreasing for each $t\geq 0$. Note that (18) implies
$|x(t)|$ $\leq\beta\left(\left\lvert\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}\right\rvert,t\right)+\varrho(|u|_{\infty})$ (19a)
$|z(t)|$ $\leq\beta\left(\left\lvert\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}\right\rvert,t\right)+\varrho(|u|_{\infty})+|h(x(t))|.$ (19b)
@@
$\beta\left(\left\lvert\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}\right\rvert,t\right)=\beta\left(\left\lvert\begin{bmatrix}x_{0}\\ z_{0}-h(x_{0})\end{bmatrix}\right\rvert,t\right)\leq\overline{\beta}\left(\left\lvert\begin{bmatrix}x_{0}\\ z_{0}\end{bmatrix}\right\rvert,t\right),$
where $\overline{\beta}(r,t):=\beta(2r,t)+\beta(2\tilde{\alpha}(r),t)\in\mathcal{GKL}$. Moreover,
$|h(x(t))|\leq\tilde{\alpha}(|x(t)|)\leq\tilde{\beta}\left(\left\lvert\begin{bmatrix}x_{0}\\ z_{0}\end{bmatrix}\right\rvert,t\right)+\tilde{\varrho}(|u|_{\infty}),$ (20)
where $\tilde{\beta}(r,t):=\tilde{\alpha}\left(2\overline{\beta}\left(r,t\right)\right)\in\mathcal{GKL}$, and $\tilde{\varrho}(\cdot):=\tilde{\alpha}(2{\varrho}(\cdot))\in\mathcal{K}$. Combining (19) and (20) yields the result.
@@ 3.3 A Stylized Example
To illustrate our results, we first consider a singularly perturbed system with scalar reduced and boundary layer systems. In particular, consider the plant
@@
where $\lceil\cdot\rfloor^{q}:=|\cdot|^{q}\text{sgn}(\cdot)$, $x,z,u_{1},u_{2}\in\mathbb{R}$, $0<q_{1}\leq r_{1}<1$ and $1<r_{2}\leq q_{2}$. System (21) has the quasi steady state $h(x)=x$, with reduced system:
@@
and boundary layer system